If we assume that a given bus market is in competition which charges a flat fare of N$1, and if the formula for the total demand (in thousands) in the market is given by the equation:

Qd = 250-60P

Where Qd is the quantity demanded in thousands at a given price P.
If we further assume constant returns to scale, then:

a) What is the total demand at the N$1 flat fare?
b) If the market is shared equally by 4 firms, what is the number of passengers per vehicle carried by each company?
c) If the cost per vehicle kilometer is N$1.60, average utilization 20 passengers per vehicle kilometer and average trip distance 10 kilometers:

i) What is the level of bus kilometers required to service this market?
ii) What profits are being made?
iii) What type of profit is this, normal or abnormal?
iv) What is the cost per passenger carried (as opposed to the cost per vehicle kilometer)?
1
Expert's answer
2015-03-31T10:52:30-0400
P = N$1, Qd = 250-60P
a) The total demand at the N$1 flat fare is: Qd = 250 - 60*1 = 190 thousands.
b) If the market is shared equally by 4 firms, the number of passengers per vehicle carried by each company is Q = 190/4 = 47.5 thousands.
c) If the cost per vehicle kilometer is N$1.60, average utilization 20 passengers per vehicle kilometer and average trip distance 10 kilometers:
i) the level of bus kilometers required to service this market is 190/20*10 = 85 thousands km.
ii) the profits are TP = TR - TC = 1*190 - 1.6*85 = N$54 thousands.
iii) the profit of N$54 thousands is abnormal.
iv) the cost per passenger carried (as opposed to the cost per vehicle kilometer) is ATC = 136/190 = N$0.72.